elsa boards sells snowboards , xpert

6-4 Elsas Boards sells a snowboard, Xpert, that is popular with snowboard enthusiasts. Information relating to Els shown below. During the same month, 264 Xpert snowboards were sold. Elsas uses a periodic inventory syste Date Explanation Units Unit Cost Total cost 99$ 5,643 10,192 4,752 11,445 $32,032 Sept. 1Inventony Sept. 12Purchases Sept. 19 Purchases Sept. 26 Purchases109 57 104 108 105 98 Totals 308 Compute the ending inventory at September 30 and cost of goods sold using the FIFO and LIFO methods. FIFO LIFO The ending inventory at September 30 s4620 Cost of goods sold INK TO TEXT 0 to search
Answer #1
FIFO:
Ending inventory at September 30 4620 =44*105
Cost of goods sold 27412 =32032-4620
LIFO:
Ending inventory at September 30 4356 =44*99
Cost of goods sold 27676 =32032-4356